OpenStudy (anonymous):

{{I will give medal}}Given that 3h+2x=2f-gx, express x in terms of f,g and h.

OpenStudy (anonymous):

@melissaaaa by mistake i closed that ques

OpenStudy (anonymous):

3h + 2x = 2f - gx subtract 3h from both sides: 2x = 2f - gx - 3h add gx to both sides: 2x + gx = 2f - 3h factor out the x: x(2+g) = 2f - 3h divide both sides by (2+g) x = (2f-3h)/(2+g)

OpenStudy (anonymous):

lol you keep closing the question ><

OpenStudy (anonymous):

hello! first get x alone on one side of the equation.

OpenStudy (anonymous):

thank yoouuu(: